Reduce the quadratic form


2 2 2 8 7 3 12 – 8 4 x y z xy yz zx    


to the canonical form


through an orthogonal transformation and hence show that it


is positive Semi-definite.
1
Expert's answer
2021-11-07T16:55:15-0500

Assuming typo error, and modifying the condition of the question into

"Reduce the quadratic form "8x^2+7y^2+3z^2-12xy-8yz+4zx"

to the canonical form through orthogonal transformation and hence show it is positive semi-definate"


Solution

Converting to matrix form


A= "\begin{bmatrix}
 8&-6& 2 \\
 -6&7& -4\\
2&-4&3
\end{bmatrix}"



Characteristic equation

det "||A-\lambda\Iota||=0"


det "\begin{Vmatrix}
 8-\lambda&-6& 2\\
 -6&7-\lambda & -4\\
2&-4&3-\lambda
\end{Vmatrix}" =0...(i)


Characteristic polynomial

"\lambda^3-D_1\lambda^2+D_2\lambda-D_3=0"

Where "D_1=" sum of main diagonal element

"=8+7+3=18"

"D_2=" sum of minors of the main diagonal element


"D_2=\begin{vmatrix}
 7& -4\\
 -4& 3
\end{vmatrix}" + "\begin{vmatrix}
 8 & 2 \\
 2 & 3
\end{vmatrix}" +


"\begin{vmatrix}
 8 & -6\\
 -6& 7
\end{vmatrix}" "=5+20+20=45"


"D_3=" Det "|A|" = "8(21-16)-"

-"6(-18+8)+2(56-36)" =


Characteristic equation is

"\lambda^3-18\lambda^2+45\lambda=0"

Solving for "\lambda;"


"\lambda=0, \>\lambda=3,\>\lambda=15"


Finding Eigenvector

Case 1 "\lambda=0"

"rref \>\begin{pmatrix}
 8&-6 & 2\\
 -6&7 & -4\\
2&-4&3
\end{pmatrix}" =


"\begin{pmatrix}
 1&0& \frac{1}{2} \\
 0&1 & -1\\
0&0&0
\end{pmatrix}"


"\begin{pmatrix}
 1&0 & \frac{-1}{2}\\
 0&1 & -1\\
0&0&0
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}" =


"\implies\>x=\frac{1}{2}z"

"y=z"


Vector ="\begin{pmatrix}
 \frac{1}{2} \\
 1 \\
1
\end{pmatrix}" ="\begin{pmatrix}
 1 \\
 2\\
2
\end{pmatrix}"


Case 2 "\lambda =3," substituting "\lambda=3" in.....(i)



"\begin{pmatrix}
 5&-6 & 2 \\
 -6&4 & -4\\
2&-4&0
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}" =0


Using ref of the matrix


"\begin{pmatrix}
 1&0& 1\\
 0&1&\frac{1}{2} \\
0&0&0
\end{pmatrix}" "\begin{pmatrix}
 x\\
 y \\
z
\end{pmatrix}" =0


"\implies\>x=-z"

"y=\>\frac{-1}{2}z" let "z=1"


Vector is "\begin{pmatrix}
 -1 \\
 \frac{-1}{2} \\
1
\end{pmatrix}" ="\begin{pmatrix}
 -2 \\
 -1\\
2
\end{pmatrix}"


Case 3, "\lambda=15," substituting "\lambda=15" in .....(i)


"\begin{pmatrix}
 -7&-6 & 2 \\
 -6&-8& -4\\
2&-4&-12
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}" =0,using rref form


"\begin{pmatrix}
 1&0& -2 \\
 0&1&2 \\
0&0&0
\end{pmatrix}\begin{pmatrix}
 x\\
 y \\
z
\end{pmatrix}" =0

"\implies \>x=2z"

"y= -2z" let "z=1"

Vector is "\begin{pmatrix}
 2 \\
 -2 \\
1
\end{pmatrix}"


Orthonormal matrix from unit eigenvectors



"Q=\frac{1}{3}" "\begin{pmatrix}
 1&-2& 2 \\
 2&-1 & -2\\
2&2&1
\end{pmatrix}"

Diagonalizing matrix


"D=Q^TAQ="

"\frac{1}{9}" "\begin{pmatrix}
 1&2 & 2 \\
 -2&-1 & 2\\
2&-2&1
\end{pmatrix}\begin{pmatrix}
 8&-6 & 2 \\
 -6&7 & -4\\
2&-4&3
\end{pmatrix}\begin{pmatrix}
 1&-2 & 2 \\
 2&-1 & -2\\
2&2&1
\end{pmatrix}"


="\begin{pmatrix}
 0&0 & 0\\
 0&3 & 0\\
0&0&15
\end{pmatrix}"


Orthogonal transformation reduce the quadratic form to canonical form


"3y_2^2+15y_3^2"


Nature of quadratic form


Contain a zero, eigenvalue and 2 eigenvalue ">1"


Therefore it is positive semi-definate